OpenStudy (anonymous):

good evening. complete each ordered pairs so that they are solutions go the given linear equation. 4x-+y=29 (9, ___), (1,____)

OpenStudy (anonymous):

4x-y=29 +y to both sides to get 4x = 29+y -29 to both side to get 4x-29 = y y=4x-29 solve f(1) solve f(9) for your answers

OpenStudy (anonymous):

I.E plug 1 in for x and solve plug 9 in for x and solve